An account is opened with an initial deposit of $6500 and earns 3.6% interest compounded semi-annually. What will the account be worth in 20 years?
Answer by jim_thompson5910(35256)

You'll use the formula A = P(1+r/n)^(n*t) where


P = 6500 is the amount deposited
r = 0.036 is the decimal form of 3.6% interest rate
n = 2 is the number of times the money is compounded per year
t = 20 is the number of years


A = P(1+r/n)^(n*t)
A = 6500(1+0.036/2)^(2*20)
A = 13,268.580898361
A = 13,268.58


So you'll have $13,268.58 in the account.